OpenStudy (anonymous):

what are the zeros of the function f(x)=6x^2-4x-2?

OpenStudy (anonymous):

start with \(6x^2-4x-2=0\) or maybe \(3x^2-2x-1=0\) after you divide by \(2\)

OpenStudy (anonymous):

factor as \((x-1)(3x+1)=0\) and then it should be ok